# ROODI: Reconstructing Occluded Objects with Denoising Inpainters

### v1

@@ 3.3 Training with occlusion-reasoned inpainting
Occlusion reasoning.
Specifically, to identify occluded areas, we look into the depth maps, one rendered using the object-centric 3D Gaussians and the other using the full scene 3D Gaussians. We then compare the two depth maps and mark pixels where the full scene is closer to the camera than the target object as to-be-inpainted pixels. We therefore define the mask $\mathbf{m}$ denoting occluded areas as:
$\mathbf{m}=\mathbbm{1}(\mathbf{d}_{\text{scene}}<\mathbf{d}_{\text{object}}),$ (5)
where $\mathbbm{1}$ again is the indicator function applied elementwise, and $\mathbf{d}_{\text{scene}}\in^{H\times W}$ and $\mathbf{d}_{\text{object}}\in^{H\times W}$ are the rendered depth maps of the full scene and the target object, respectively. As this mask can sometimes be unreliable due to reconstruction errors, we relax the mask to account for errors via morphological opening as shown in Fig. 4 with a $17\times 17$ kernel. Note that the mask, shown as solid white, also contains the background. This allows the inpainter to also clean up the borders of the object.
@@
Inpainting.
To inpaint occluded areas, we generate four random camera views for each object, with the object centered and captured from various angles. To maintain multi-view consistency, following [51], we arrange these four renderings into a $2\times 2$ grid view and provide them to the inpainter simultaneously. With the occlusion masks for each view from Eq. 5, we use an off-the-shelf inpainting diffusion model, “SD-XL Inpainting 0.1”  [46], which is modified and finetuned from SDXL [33]. We do not use any prompts for the inpainter. The inpainting outcome, with our mask, provides enhancements as shown in Fig. 4. We aggregate these enhancements into the object-centric Gaussians via training, which we detail next.
@@
Figure 5: Example views from MultiObjectBlender – We synthesize four scenes with ShapeNet [5] assets. In each scene we select three to four objects, each marked with red arrow, as the objects of interest. We provide stand-alone renders for each objects as test images.
Training with inpainted images.
We then further train the object-centric 3DGS with the resulting four inpainted views. Similar to the original 3DGS training, we rely on the $\mathcal{L}_{1}$ loss to guide each pixel. However, we do not use the D-SSIM loss as we find it to be harmful as inpainting results are not always perfect. Instead, we use a perceptual loss to guide the inpainting process as in Wu et al. [53]:
where $\lambda_{L_{1}-Inpaint}$ and $\lambda_{Perc}$ are hyperparameters balancing the two losses, and $\mathcal{L}_{Perceptual}$ is the perceptual distance [64]. We finetune the object-centric 3DGS for a total of 100 iterations, using the original image (the full scene) every iteration and the inpainted images every other iteration, to prevent inpainting from causing our reconstructions to drift away too much from the original object.
4 Experiments
Original Image Gaussian Grouping SAGA OmniSeg3D-GS Our Method
@@ 4 Experiments
figurines (zoomed) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
ramen (zoomed) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
teatime Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative examples from the LERF [19] dataset – We show example renders, with the selected objects marked with red arrows. We show the same view for all methods, with an extra view for our method. Our method provides the best extractions, even for occluded objects and in challenging scenarios.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Bedroom Kitchen Livingroom Picnic Average
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
OmniSeg3D-GS [59] 36.34 0.7877 0.1924 30.12 0.5789 0.3526 34.61 0.8575 0.1274 36.82 0.9108 0.0849 34.47 0.7837 0.1893
Gaussian Grouping [57] 31.82 0.7189 0.2708 29.98 0.5564 0.4074 31.21 0.6825 0.3020 31.68 0.7351 0.2598 31.17 0.6732 0.3100
SAGA [3] 33.06 0.7703 0.1993 33.56 0.7921 0.1777 29.79 0.4566 0.4808 34.69 0.9016 0.0971 32.78 0.7302 0.2387
Ours 37.62 0.9402 0.0710 34.54 0.8733 0.1238 35.05 0.9109 0.0860 36.47 0.9069 0.0964 35.87 0.9084 0.0934
Table 1: Quantitative summary for MultiObjectBlender – We report quantitative summary over all objects for each scene, as well as the average. We mark best scores with bold and second-bset with an underline. Our method performs either best or second-best on all metrics, and significantly outperforms all methods on average.
@@
Figure 8: Ablation study – (top row) Numerous floaters remain without pruning, and pruning alone (w/o inpainting) removes most but leaves some near the spoon. Our method eliminates both distant and nearby floaters via pruning and inpainting, respectively, resulting in clean edges. (bottom row) Without inpainting, occluded objects exhibit holes. Both non-inpainting enhancement and inpainting fill these holes, but non-inpainting enhancement also edits other correct regions as well, drifting away from the actual object appearance.
@@ 4.1 Experimental setup
We evaluate our method on two datasets: the LERF dataset [19], a commonly used real-world dataset for extraction tasks; and on our custom synthetic dataset, which we name MultiObjectBlender.
LERF [19] dataset.
@@
The LERF dataset contains 14 scenes, among which we evaluate our method on three scenes as in prior work [57, 4, 9]—we use figurines, ramen, and teatime. Being a real-world dataset, the ground-truth views for a single object are non-trivial to obtain because of occlusion. A strightforward option is to take the masks of the object of interest on the ground-truth images and evaluate within the masks, but this would ignore the quality of reconstructions outside the masks—e.g., the floaters. Due to this, existing works [57, 9] evaluate rendered semantic features against the semantic segmentation masks, but this is not applicable to our case as we also reconstruct occluded regions. We thus use this dataset only for qualitative evaluation, and rely on our synthetic dataset for quantitative evaluation.
MultiObjectBlender dataset.
@@
To quantitatively evaluate the quality of novel-view synthesis, even for occluded regions, we create a new synthetic dataset, MultiObjectBlender. We use the ShapeNet [5] assets and render four complex scenes as shown in Fig. 5. Each scene contains 5 to 20 objects, and we select 3 to 4 objects per scene for testing marked with red arrows. Each scene has distinct characteristics, such as a large number of highly reflective objects, strong shadows from sunlight in outdoor settings, or significant occlusions where objects overlap. To allow quantitative evaluation, we render both the full scene and the object of interest in isolation, without other objects or backgrounds. We render 50 training images of the full scene and 50 test images for each extraction target object.
Evaluation metrics
We evaluate the quality of extraction using three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Metric (SSIM) [50], and Learned Perceptual Image Patch Similarity (LPIPS) [64]. As we wish to focus on the object of interest, we only compute these metrics for the pixels within a bounding box defined by 1.2$\times$ the bounding box that tightly bounds the target object. We average the metrics across all extracted objects for each scene.
Baselines
We compare our method with three baselines:
@@
  • OmniSeg3D-GS [59] (CVPR’24): is an author-provided extension of OmniSeg3D [59] that uses 3DGS as the underlying 3D representation. It trains 3DGS with an additional feature that can be used to segment out objects. We select a pixel on the target object and extract Gaussians whose cosine similarity of semantic feature with the target object exceeds a preset threshold. We carefully choose the threshold and the pixel for best results.
@@
  • Gaussian Grouping [57] (ECCV’24): trains 3DGS with an additional feature, and with a classifier that takes each feature as input and outputs logits. We extract Gaussians where the logit corresponding to the target is maximized.
@@
  • SAGA [3] (AAAI’25): trains 3DGS with an additional feature, similar to the former two. As in the case of OmniSeg3D-GS, we again select a pixel on the target object and extract Gaussians whose cosine similarity exceeds a preset threshold.
@@ 4.2 Results
Qualitative results – Figs. 6 and 7.
We first present qualitative results. We show rendered images of extracted objects on the LERF dataset in Fig. 6 and on the MultiObjectBlender dataset in Fig. 7. In both datasets, our method provides improved renderings, with less floaters and holes. Note especially the top of the cube in Fig. 6 which is occluded in all views, or the book in Fig. 7 where other methods struggle to reconstruct.
Quantitative results – Tab. 1.
We provide quantitative results for the MultiObjectBlender dataset in Tab. 1. As reported, our method significantly outperforms all baselines. Notably, the gap is larger when looking into SSIM and LPIPS, which reflects the quality difference shown in Figs. 6 and 7. In the case of PSNR, the gap is relatively smaller as our metric still contains fully white backgrounds to account for also floaters, which are prevalent when trying to extract objects in cluttered scenes.
4.3 Ablation studies – Fig. 8.
To motivate our design choices, we perform ablation studies. We show importance of pruning based on K-th nearest neighbor and inpainting qualitatively in Fig. 8. As shown, our method without pruning results in many floaters. Our method without inpainting is also suboptimal, as nearby floaters remaining after pruning cause unclear object edges. The two together provides the best results.
We also show that using a non-inpainting diffusion model to enhance the quality of rendering also does not drastically improve results. Specifically, instead of inpainting, we use text-to-image denoising model “SD-XL 1.0-base” [33] with the SDEdit algorithm [28] to enhance the quality of renderings. This, however, generates enhanced images that are very inconsistent from different views due to lack of preservation of the unoccluded content. This results in quality degradation and can even drift away from how the original object is observed in some cases.
@@ 5 Conclusion
We presented a novel approach for extracting objects from 3D Gaussian Splatting models. Our method addresses two key challenges: isolating individual 3D Gaussian primitives for each object and handling occlusions in complex scenes. By implementing K-nearest neighbor pruning to remove irrelevant primitives and leveraging diffusion-based inpainting to compensate for occluded regions, our approach significantly outperforms existing methods. We demonstrated the synergy between pruning and inpainting, both of which substantially enhance extraction performance. Our evaluations on both the standard LERF dataset and our new synthetic MultiObjectBlender  dataset confirm the effectiveness of our approach in producing clear object extractions that are robust to occlusions and floaters.
Limitations.
While our method shows promising results, the quality currently is limited by the generative inpainting model. As research in this area is rapidly evolving, we expect this limitation to be mitigated in the future.

### v2

@@ 3.3 Training with occlusion-reasoned inpainting
Occlusion reasoning. Specifically, to identify occluded areas, we look into the depth maps, one rendered using the object-centric 3D Gaussians and the other using the full scene 3D Gaussians. We then compare the two depth maps and mark pixels where the full scene is closer to the camera than the target object as to-be-inpainted pixels. We therefore define the mask $\mathbf{m}$ denoting occluded areas as:
$\mathbf{m}=\mathds{1}(\mathbf{d}_{\text{scene}}<\mathbf{d}_{\text{object}}),$ (6)
where $\mathds{1}$ again is the indicator function applied elementwise, and $\mathbf{d}_{\text{scene}}\in\mathds{R}^{H\times W}$ and $\mathbf{d}_{\text{object}}\in\mathds{R}^{H\times W}$ are the rendered depth maps of the full scene and the target object, respectively. As this mask can sometimes be unreliable due to reconstruction errors, we relax the mask to account for errors via morphological opening as shown in Fig. 3 with a $17\times 17$ kernel. Note that the mask, shown as solid white, also contains the background. This allows the inpainter to also clean up the borders of the object.
@@
Inpainting. To inpaint occluded areas, we generate four random camera views for each object, with the object centered. To maintain multi-view consistency, following NeRFiller (Weber et al. 2024), we arrange these four renderings into a $2\times 2$ grid view and provide them to the inpainter simultaneously. With the occlusion masks for each view from Eq. 6, we use an off-the-shelf inpainting diffusion model, “SD-XL Inpainting 0.1”  (2023), which is modified and finetuned from SDXL (Podell et al. 2023). We do not use any prompts for the inpainter. The inpainting outcome, with our mask, provides enhancements as shown in Fig. 3. We aggregate these enhancements into the object-centric Gaussians via training, which we detail next.
Training with inpainted images. We then further train the object-centric 3DGS with the resulting four inpainted views. Similar to the original 3DGS training, we rely on the $\mathcal{L}_{1}$ loss to guide each pixel. However, we do not use the D-SSIM loss as we find it to be harmful as inpainting results are not always perfect. Instead, we use a perceptual loss to guide the inpainting process as in ReconFusion (Wu et al. 2024b):
where $\lambda_{L_{1}-Inpaint}$ and $\lambda_{Perc}$ are hyperparameters balancing the two losses, and $\mathcal{L}_{Perceptual}$ is the perceptual distance (Zhang et al. 2018). We finetune the object-centric 3DGS for a total of 100 iterations, using the original image (the full scene) every iteration and the inpainted images every other iteration, to prevent inpainting from causing our reconstructions to drift away too much from the original object.
4 Experiments
@@
Figure 4: Example views from MultiObjectBlender – We synthesize four scenes with ShapeNet (Chang et al. 2015) assets. In each scene we select three to four objects, each marked with red arrow, as the objects of interest. We provide stand-alone renders for each objects as test images.
Original Image Original (zoomed) Gaussian Grouping SAGA OmniSeg3D-GS Our Method
@@
teatime Refer to caption N/A Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative examples from the LERF dataset – We show example renders, with the selected objects marked with red arrows. Our method provides the best extractions, even for occluded objects and in challenging scenarios.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ IoU $\uparrow$
OmniSeg3D-GS (2024) 34.47 0.7837 0.1893 0.5388
Gaussian Grouping (2024) 31.17 0.6732 0.3100 0.3012
SAGA (2025) 32.78 0.7302 0.2387 0.5664
Ours 36.12 0.9164 0.0802 0.9419
Table 1: Quantitative summary for MultiObjectBlender – We report quantitative summary over all objects for each scene, as well as the average. Our method significantly outperforms all methods.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ IoU $\uparrow$
w/o-two-stage 30.59 0.6538 0.3501 0.3031
w/o-inpainting 36.02 0.9193 0.0755 0.9418
w/o-pruning 35.94 0.9116 0.0874 0.9255
w/o-Wasserstein (Euclidean) 36.05 0.9144 0.0816 0.9372
Ours 36.12 0.9164 0.0802 0.9419
Table 2: Quantitative ablation for MultiObjectBlender We report ablation with various components disabled. Our full method performs best.
@@
Figure 7: Ablation study – (top row) Numerous floaters remain without pruning. Pruning alone (w/o inpainting) still leaves floaters near the pot. Pruning and inpainting together results in clean edges. (bottom row) Without inpainting, occluded objects exhibit holes. Both non-inpainting enhancement and inpainting fill these holes, but non-inpainting enhancement also edits other correct regions, drifting away from the actual object appearance.
@@ 4.1 Experimental setup
We evaluate our method on two datasets: the LERF dataset (Kerr et al. 2023), a commonly used real-world dataset for extraction tasks; and on our custom synthetic dataset, which we name MultiObjectBlender.
@@
LERF (Kerr et al. 2023) dataset. The LERF dataset contains 14 scenes, among which we evaluate our method on three scenes as in prior work (Ye et al. 2024; Chacko et al. 2025; Choi et al. 2024)—we use figurines, ramen, and teatime. Being a real-world dataset, the ground-truth views for a single object are non-trivial to obtain because of occlusion. A strightforward option is to take the masks of the object of interest on the ground-truth images and evaluate within the masks, but this would ignore the quality of reconstructions outside the masks—e.g., the floaters. Due to this, existing works (Ye et al. 2024; Choi et al. 2024) evaluate rendered semantic features against the semantic segmentation masks, but this is not applicable to our case as we also reconstruct occluded regions. We thus use this dataset only for qualitative evaluation, and rely on our synthetic dataset for quantitative evaluation.
@@
MultiObjectBlender dataset. To quantitatively evaluate the quality of novel-view synthesis, even for occluded regions, we create a new synthetic dataset, MultiObjectBlender. We use the ShapeNet (Chang et al. 2015) assets and render four complex scenes as shown in fig:multiobjectblender. Each scene contains 5 to 20 objects, and we select 3 to 4 objects per scene for testing marked with red arrows. Each scene has distinct characteristics, such as a large number of highly reflective objects, strong shadows from sunlight in outdoor settings, or significant occlusions where objects overlap. To allow quantitative evaluation, we render both the full scene and the object of interest in isolation, without other objects or backgrounds. We render 50 training images of the full scene and 50 test images for each extraction target object.
[…]
Baselines We compare our method with three baselines:
@@
  • OmniSeg3D-GS (Ying et al. 2024): is an author-provided extension of OmniSeg3D (Ying et al. 2024) that uses 3DGS as the 3D representation. It trains 3DGS with an additional feature that can be used to segment objects. We select a pixel on the target object and extract Gaussians whose cosine similarity of semantic feature with the target object exceeds a preset threshold. We carefully choose the threshold and the pixel for best results.
@@
  • Gaussian Grouping (Ye et al. 2024): trains 3DGS with an additional feature, and with a classifier that takes each feature as input and outputs logits. We extract Gaussians where the logit corresponding to the target is maximized.
@@
  • SAGA (Cen et al. 2025): trains 3DGS with an additional feature, similar to the former two. As in the case of OmniSeg3D-GS, we again select a pixel on the target object and extract Gaussians whose cosine similarity exceeds a preset threshold.
@@ 4.2 Results
Qualitative results – Figs. 5 and 6. We first present qualitative results. We show rendered images of extracted objects on the LERF dataset in Fig. 5 and on the MultiObjectBlender dataset in Fig. 6. In both datasets, our method provides improved renderings, with less floaters and holes. Note especially the top of the cube and the lower half of the egg in Fig. 5, which is occluded in all views, or the book in Fig. 6, where other methods struggle to reconstruct.
Quantitative results – Table 1. We provide quantitative results for the MultiObjectBlender dataset in Table 1. As reported, our method significantly outperforms all baselines. Notably, the gap is larger when looking into SSIM and LPIPS, which reflects the quality difference shown in Figs. 5 and 6. In the case of PSNR, the gap is relatively smaller as our metric still contains fully white backgrounds to account for also floaters, which are prevalent when trying to extract objects in cluttered scenes.
4.3 Ablation studies – Table 2, Fig. 7.
To motivate our design choices, we perform ablation studies. We show the importance of our statistical pruning and inpainting qualitatively in Fig. 7. As shown, our method without pruning results in many floaters. Our method without inpainting is also suboptimal, as nearby floaters remain after pruning, causing unclear object edges. The two together provide the best results. Table 2 quantifies this; the complete pipeline shows the best results for PSNR and IoU, and second-best for SSIM and LPIPS, where w/o-inpainting attains a marginally higher score due to its sensitivity to minor chromatic shifts.
We also show that using a non-inpainting diffusion model to enhance the quality of rendering also does not drastically improve results. Specifically, instead of inpainting, we use text-to-image denoising model “SD-XL 1.0-base” (Podell et al. 2023) with the SDEdit algorithm (Meng et al. 2021) to enhance the quality of renderings. This, however, generates enhanced images that are very inconsistent from different views due to lack of preservation of the unoccluded content. This results in quality degradation and can even drift away from how the original object is observed in some cases.
@@ 5 Conclusion
We presented a novel approach for extracting objects from 3D Gaussian Splatting models. To do so, we have proposed a prune-and-inpaint method that applies statistical outlier removal of non-object floaters, and a scene-geometry-based occlusion reasoning method for inpainting. We demonstrated the synergy between pruning and inpainting, both of which substantially enhance extraction performance. Our evaluations on both the standard LERF dataset and our new synthetic MultiObjectBlender  dataset confirm the effectiveness of our approach in producing clear object extractions that are robust to occlusions and floaters.
Limitations. While our method shows promising results, the quality currently is limited by the generative inpainting model. As research in this area is rapidly evolving, we expect this limitation to be mitigated in the future.